OpenStudy (ny,ny):

find slope and y intercept: 2y+4x=0

OpenStudy (campbell_st):

rewrite the equation in point slope form y=mx + b steps 1 subtract 4x from both sides 2. divide by 2

OpenStudy (ny,ny):

oh yeahh!!!

OpenStudy (ny,ny):

wait i got this

OpenStudy (ny,ny):

2y+4x=0 2y=0-4x y=0-4x/2

OpenStudy (campbell_st):

yes... you don't need to worry about zero and simplify the fraction y = -2x

OpenStudy (ny,ny):

slope = -2/1 y intercept = 0 right?

OpenStudy (ny,ny):

y intercept = (0,0)

OpenStudy (campbell_st):

yes... but again... write the slope as -2 and not -2/1

OpenStudy (ny,ny):

ya, my teacher says we gotta put it over one

OpenStudy (ny,ny):

but same thing

OpenStudy (campbell_st):

ok.... If thats what they say... good job

OpenStudy (ny,ny):

becuz it helps when we graph

OpenStudy (campbell_st):

ok...

OpenStudy (ny,ny):

anyway, thank yu!!!! =)